OpenStudy (anonymous):

A card is drawn at random from a deck of 52 playing cards. Find the probability of drawing an ace or a 7

OpenStudy (therealmeeeee):

In a deck there a 52 cards with 4 suits(hearts,spades,diamonds and clovers). 2 suits are red while the other 2 are black. And there are 4 cards for each number.(ace,jack.queen etc.)And remember an ace isnt considered a face card. 1) there are 2 red kings in any given deck so probability is 2/52=1/26 2)since there are 4 cards of each you multiply 4*3=12 probability is 12/52=3/13 3)I have no idea what a fade is. 4)Remember that there are 4 suits? Each suit has 13 cards. So as there are 2 suits probability is 13*2/52=26/52=1/2 5)There are 26 black cards,and they contain 6 face cards. And there are 26 red cards which contain only 2 red aces. so add both 8/52=2/13 6)In a deck there are 12 face cards (king, queen, and jack) and 13 cards of spades 3 face cards are already in spades so subtract 3. (9+13)/52=22/52=11/26

OpenStudy (therealmeeeee):

do this help

OpenStudy (anonymous):

Yes you are so helpful! Thank you!

OpenStudy (therealmeeeee):

click best response

OpenStudy (therealmeeeee):

ty

OpenStudy (anonymous):

yes master